OpenStudy (anonymous):

Use trial and error to find 2 by 2 examples of -Matrices A and B with AB = −BA but AB not equal to 0

OpenStudy (anonymous):

This is a trial and error problem. So you just need to plug in numbers. For easy calculations I would only use 0,1,-1

OpenStudy (anonymous):

Try this for example \[\left[\begin{matrix}1 & 0 \\ 1 &1\end{matrix}\right]\]

OpenStudy (anonymous):

The above is A And for B: \[\left[\begin{matrix}1 & -1 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

ok i think i know how to do it much appreciate